OpenStudy (anonymous):

A lottary has a million dollar first prizes, a twenty five thousand dollar second prizes, five of thousand dollar third prizes. If a ticket cost 2 dollar, what is the expected profit per ticket?

OpenStudy (valpey):

I'm sure you mean "has a million dollar first prize,..." Expected profit will be $1M * Probability of winning first + $25K * Probability of winning second + $1K * Probability of winning third - $2 (price of ticket). Usually, the answer will be negative. We need to know something about the chance of winning.

OpenStudy (anonymous):

chance of winning is 7/2million

OpenStudy (valpey):

So the chance of winning first is 1:2,000,000, the chance of winning second is 1:2,000,000, and the chance of winning third is 5:2,000,000?

OpenStudy (valpey):

If so, then Expected Profit is:\[\frac{$1,000,000}{2,000,000}+\frac{$25,000}{2,000,000}+\frac{5*$1,000}{2,000,000}-$2\]

OpenStudy (anonymous):

Yey! its right but it should be positive so I think we can sutract the ratio from $2 Many many Thanx